Over four years the price of a car decreased from $40000 by 55%. What is the price of the car now?
Answer by stanbon(75887)

Over four years the price of a car decreased from $40000 by 55%. What is the price of the car now?
-------
40000-0.55(40000) = 0.45*40000 = $18,000
-------------------------------
Cheers,
Stan H.
==============